The management of Diversified Products Limited (DPL) are investigating whether to increase plant capacity. DPL is a manufacturing company that produces an array of products. One of its products, the Pod, has a selling price of $80 per unit. Based on a market survey, which cost $68 500, the marketing director is convinced that sales will increase for the next five years if the company was to increase current capacity. The marketing manager believes that capacity will increase to 35 000 should DPL buy a new plant. If DPL leased a new plant, there would be extra efficiencies and the capacity would increase to 40 000. The marketing manager will earn a bonus of $90 000 at the end of year three if his estimations are correct. The company’s current capacity is 30 000 units per annum. Each unit has a variable cost of $40 and an allocated fixed cost of $15, based on a capacity of 30 000 units per annum. A new plant could be purchased for $850 000. The new plant has an estimated useful life of five years, at the end of which the salvage value is expected to be $230 000. It is estimated that because of technological advances, the new plant will result in increased efficiencies and a reduction in variable costs of $10 per unit. Fixed costs, other than depreciation, are expected to increase by $5 per unit. The wear and tear allowance for the new machine is the same as the accounting depreciation. The lease payments would be $220 000, payable annually in advance for five years. The company’s WACC is 15%, the cost of equity 18% and the debt-to-equity ratio is 0.6. The company’s tax rate is 28% and taxation is paid annually at the end of the year. Required: Advise management in a letter as to whether they should buy or lease the new plant.
Answer #1

Selling Price of the Pod = 80 $

The cost of the market survey is irrespective of which option the company chooses ( buy or lease) and is a sunk cost and hence does not impact the buy or lease decision.

Current Capacity = 30,000 units per annum , Present variable cost per unit = 40 $ , Allocated fixed cost per unit = 15 $

a) Buy Option

Cost of the new plant = 8,50,000 $ , Useful life = 5 years , Salvage Value = 230,000 $

Assuming a straight line depreciation the annual depreciation is = (8,50,000 - 2,30,000)/ 5 =6,20,000/5 = 1,24,000 $

Variable cost with the new plant = 40 - 10 = 30 $ , Allocated fixed cost per unit = 15 + 5 = 20 $

Capacity of the new plant = 35,000 , Addition is capacity = 35000 - 30000 = 5000 units

Incrementa cash flow with the buy option

Reduction in cost for 30,000 units from 55 $ per unit to 50 $ per unit which is a saving of 5 $ per unit

For 30,000 units this is an incremental cash flow of 30,000 * 5 = 1,50,000 $

The profit margin from the additional 5000 units = (80 - 50) * 5000 = 30 * 5000 = 1,50,000 $

Total additional cash flow = 1,50,000 + 1,50,000 - 1,24,000 = 3,00,000 - 1,24,000 = 1,76,000 $

So the company will get this additional cashflow for 5 years if it buys the new plant. The company's wacc is 15% which is the opportunity cost for this.

So using this the NPV of the buy option is = NPV(.15,1,76,000,1,76,000,1,76,000,1,76,000,1,76,000) = 589,979 $

b) Lease option

With lease option the capacity will increase to 40,000 unit which is an increase of 40,000 - 30,000 = 10,000 units

The profit margin from the additional 10,000 units = (80-50) * 10,000 = 30 * 10,000 = 3,00,000 $

The incremental cash flow with lease option = 1,50,000 + 3,00,000 = 4,50,000 $ per year

The NPV of this cash flow is = NPV(.15,4,50,000,4,50,000,4,50,000,4,50,000,4,50,000) = 1,508,469 $

However the company needs to pay lease rental which is a cash outflow.

Annual lease rental = 2,20,000 $ , tax rate = 28%

Since lease rentals are tax deductible the company will pay an effective lease rental of = 2,20,000 * ( 1- .28)

= 1,58,400

Also the lease rentals need to be discounted at company's cost of debt which can be computed from wacc.

.15 = .625 * .18 + .375 ^ Rd * .72

.0375 = .375 * Rd * .72

Rd = 13.88 %

NPV of the lease rental discounted at cost of debt which is paid at the beginning of year is = PV(.1388,5,-158400,0,1) = 621 069 $

This must be deducted the cash inflow as this is the net present value of cash outflow with lease rental

So the NPV of lease option = 1,508,469 - 621069 = 887,400 $

So the lease option still has a higher NPV than the buy option (589,979 $) and so the company should lease the new plant